OpenStudy (anonymous):

Henry's Law constants for aqueous solutions at 25 °C are 8.20 × 10-7 molal/mmHg for N2 and 1.62 × 10-6 molal/mmHg for O2. Determine the solubility of oxygen in water under an atmospheric pressure of 760 mmHg, assuming that air is 80% N2 and 20% O2.

OpenStudy (anonymous):

(760 * 20%)=152 152 * 1.62 * 10^-6 = 2.46 * 10^-4

OpenStudy (anonymous):

can someone tell me if I am correct?